Grade 11Mechanics

The pressure at the bottom in a lake is 3/2 times that at half depth where the barometer reads 10m . The depth of the lake is ?

Nolan Noronha
7 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Let the point at half depth be x,
P1=P+dgx
Let the depth of lake be 2x,
P2=P+dg2x
P2=3/2×P1 
2P2=3P1 
2 (P+dg2x)=3 (P+dgx)
2P+4dgx=3P + 3dgx
P=dgx
x=P/dg
x=10^5/10000 (d=10^3kg/m^3) (P=10^5pa)
But depth of lake is 2x
2x=2×10=20m